# A preorder on the set of links defined via orbifolds

### v1

@@ Abstract.
For a link $L$ in the $3$-sphere, the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ is defined as a quotient of the link group of $L$. When there exists an epimorphism $G^{\mathrm{orb}}(L)\to G^{\mathrm{orb}}(L^{\prime})$, we denote this by $L\succeq L^{\prime}$ and explore the relationships between the two links. Specifically, we prove that if $L\succeq L^{\prime}$ and $L$ is a Montesinos link with $r$ rational tangles $(r\geq 3)$, then $L^{\prime}$ is either a Montesinos link with at most $r+1$ rational tangles or a certain connected sum. We further show that if $L$ is a small link, then there are only finitely many links $L^{\prime}$ satisfying $L\succeq L^{\prime}$. In contrast, if $L$ has determinant zero, then $L\succeq L^{\prime}$ for every $2$-bridge link $L^{\prime}$. Additionally, we discuss applications to symmetric unions of knots and connections to other preorders on the set of knots. Finally, we raise open questions on bridge number and volume.
@@ 1.1. A preorder defined via orbifolds
    If $L$ is a Seifert link whose determinant is non-zero, then $L^{\prime}$ is the unknot, or a $2$-bridge link, or an elliptic Montesinos link or a Seifert link, or a connected sum of $n_{1}$ 22-bridge links and $n_{2}$ elliptic Montesinos links with $n_{1}+2n_{2}\leq\operatorname{rank}(\pi_{1}(\Sigma_{2}(L)))$.
@@ 1.2. Arborescent links
22-bridge links, Montesinos links, and Seifert links belong to a much wider class of links called arborescent links. A link $L$ is an arborescent link if $\Sigma_{2}(L)$ is a graph manifold, that is to say the geometric decomposition of $\Sigma_{2}(L)$ along essential 2-spheres and tori involves only Seifert fibered pieces. See [14], [39, Chapter 10.7], and [58]. For example, the Kinoshita-Terasaka knot is an arborescent knot. We also get some constraints on a link $\pi$-dominated by an arborescent link.
@@ 1.3. Small links and finiteness
Finally, we recall that a symmetric union construction of links gives a $\pi$-domination. The main results (Theorem 1.5 and 1.11) can be applied to the study of symmetric unions and to deduce Corollaries 7.3, 7.4, and 7.5.
@@ Acknowledgments
This study was supported in part by JSPS KAKENHI Grant Numbers JP19K03505 and JP23K12974. The first and second authors were supported by Soka University International Collaborative Research Grant.
@@ 2. Preliminaries
According to [39, Definition 3.2.2], a link $L\subset S^{3}$ is said to be locally trivial if any $2$-sphere $S$ intersecting $L$ in two points bounds a $3$-ball $B$ such that $B\cap L$ is a trivial arc in $B$. The unknot $U$ is locally trivial. For a split link, it is locally trivial if and only if it is a $2$-component trivial link $U\sqcup U$. A connected sum of two links distinct from the unknot is not locally trivial.
Definition 2.1 ([39, Definition 3.2.4]).
@@
The following factorization theorem will be useful (see [31, Main Theorem], [39, Theorem 3.2.6]).
@@ 2.1. Montesinos links
See [36, Corollary 6.2] for instance. Moreover, $r-2\leq\operatorname{rank}(\pi_{1}(\Sigma_{2}(L)))\leq r-1$ by [10, Theorem 1.1].
The Seifert fibered manifold $\Sigma_{2}(L)=V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ is determined, up to orientation-preserving homeomorphism, by the rational Euler number $e_{0}\in\mathbb{Q}$ and the set of fractions $\{\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}}\}$ in $\mathbb{Q}/\mathbb{Z}$ up to permutations (see [55], [36, Theorem 1.5]), while the Montesinos link $L(\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ is determined, up to orientation-preserving homeomorphism of $S^{3}$, by the rational Euler number $e_{0}\in\mathbb{Q}$ and the set of fractions $\{\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}}\}$ in $\mathbb{Q}/\mathbb{Z}$ up to dihedral permutations (see [17, Theorem 12.26]). Note that the sign convention for the rational Euler number $e_{0}=\sum_{i}\frac{\beta_{i}}{\alpha_{i}}$ is opposite to the choice made in [53, Section 1].
@@ 2.2. Orbifolds
(1) follows from the proof of the Smith conjecture (see [52]). (2) holds by construction. (3) is proved by the equivariant sphere theorem (see [47, Section 7] or [24]).
@@
Suppose, to the contrary, that $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is centerless. Then $\tilde{\varphi}(Z)=\{1\}$ and $\tilde{\varphi}$ induces an epimorphism from the orbifold group $\Gamma=\pi_{1}(\Sigma_{2}(L))/Z$ onto $\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $\Sigma_{2}(L^{\prime})$ is aspherical, $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is torsion-free. On the other hand, $\Gamma$ is generated by torsion elements. Thus, the image $\tilde{\varphi}(\Gamma)$ must be trivial, and this is impossible. Therefore, $\pi_{1}(\Sigma_{2}(L^{\prime}))$ has a non-trivial center. In particular, it contains an infinite cyclic normal subgroup and it follows from [18, Theorem 1.1] or [27, Corollary 2] that $\Sigma_{2}(L^{\prime})$ admits a Seifert fibration. By the orbifold theorem (see [48]), one can assume that the covering involution $\tau^{\prime}$ on $\Sigma_{2}(L^{\prime})$ is fiber-preserving. If $\tau^{\prime}$ preserves the orientation of the fibers, $L^{\prime}$ is a Seifert link. Otherwise, the underlying space of the base of the Seifert fibration is $S^{2}$ by [49, Section 5.11] and $L^{\prime}$ is a Montesinos link with $r^{\prime}$ rational tangles by [49, Section 2], where
$r^{\prime}\leq\operatorname{rank}(\pi_{1}(\Sigma_{2}(L^{\prime})))+2\leq\operatorname{rank}(\pi_{1}(\Sigma_{2}(L)))+2\leq r+1.$
The center $Z=\langle h\rangle$ of $\pi_{1}(\Sigma_{2}(L))$ is a infinite cyclic normal subgroup in $G^{\mathrm{orb}}(L)$ generated by an element $h$ such that $\bar{\mu}h\bar{\mu}^{-1}=h^{-1}$, where $\bar{\mu}$ is the image of a meridian $\mu$ of $G(L)$ in the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$. This follows from the fact that, for a Montesinos link, the covering involution $\tau$ on $\Sigma_{2}(L)$ is fiber-preserving and reverses the orientation of the fibers (see [49, Section 2] and [12, Theorem 1.3]). The image $\varphi(Z)$ of the center $Z$ of $\pi_{1}(\Sigma_{2}(L))$ is in the center of $\pi_{1}(\Sigma_{2}(L^{\prime}))$.
@@
We now assume that $L^{\prime}$ is not split. Since $L^{\prime}$ is not prime by Theorem 2.2, $L^{\prime}$ decomposes into a connected sum $L^{\prime}=L^{\prime}_{1}\sharp\cdots\sharp L^{\prime}_{n}$ of finitely many prime links which are unique up to permutations ($n\geq 2$). Then, by Lemma 3.1(1)–(3), $\Sigma_{2}(L^{\prime})\cong\Sigma_{2}(L^{\prime}_{1})\sharp\cdots\sharp\Sigma_{2}(L^{\prime}_{n})$ is a non-trivial connected sum of closed, orientable, irreducible $3$-manifolds. In particular, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\cdots\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{n}))$ is a non-trivial free product and hence centerless. The argument given above for the split case shows that each factor $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ must contain some torsion element ($1\leq i\leq n$). Since $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ is the fundamental group of a closed irreducible orientable $3$-manifold for $1\leq i\leq n$, it must be of finite order by [33, Corollary 9.9]. By the orbifold theorem, each $\Sigma_{2}(L^{\prime}_{i})$ is a Seifert $3$-manifold with finite fundamental group, and each link $L^{\prime}_{i}$ is either a $2$-bridge link or an elliptic Montesinos link with three rational tangles. Let $n_{1}$ be the number of 2-bridge factors of $L^{\prime}$ and $n_{2}$ the number of elliptic Montesinos factors of $L^{\prime}$. Then, the number $n=n_{1}+n_{2}$ of prime factors of $L^{\prime}$ verifies $n_{1}+2n_{2}\leq r-1$ since the fundamental group of the $2$-fold branched cover of a $2$-bridge link has rank $1$, that of an elliptic Montesinos link has rank $2$, and that of a Montesinos link with $r\geq 3$ rational tangles has rank at most r1r-1.
@@
By [11], the bridge number $b(K)$ of a Montesinos knot with $r\geq 3$ tangles is $r$. Then, the proof follows readily from Theorem 1.5(3) when $K^{\prime}$ is prime. When $K^{\prime}$ is not prime, it is the connected sum of $n_{1}$ 22-bridge knots and $n_{2}$ elliptic Montesinos knots with $n_{1}+2n_{2}\leq r-1$. Schubert’s bridge formula [59] shows that the bridge number minus one is additive by connected sum:
@@
Let $N$ be a closed orientable irreducible $3$-manifold whose JSJ pieces are all hyperbolic. Then $N$ is aspherical and thus $\pi_{1}(N)$ is torsion-free. Moreover, by [35, Theorem VI.1.6(i)], the centralizer in $\pi_{1}(N)$ of any non-trivial element is isomorphic to $\mathbb{Z}$ or $\mathbb{Z}\oplus\mathbb{Z}$ (see also [26]), hence abelian. Then, Theorem 1.8 follows from Theorem 4.7 and Corollary 4.10, where the proof of Theorem 4.7 is based on Lemmas 4.1 and 4.3 below.
@@
Since $M_{0}$ is a rational homology solid torus, one can choose a simple closed curve $\alpha\subset\partial M_{0}$ such that the Dehn filling $M_{0}(\alpha)$ is a rational homology sphere (see [64, Lemma 3.2] for example). The natural epimorphism
@@
Combining Theorem 4.7 with [35, Theorem VI.1.6(i)], one has the next corollary.
@@ 5. Small links and finiteness
If $L$ is not a split link, $E(L)$ is irreducible and a properly embedded essential surface $F$ in $E(L)$ is boundary-incompressible, see [54, Lemma 2.1]. If $\Sigma_{2}(L)$ does not contain any incompressible surface, then the link $L$ is small by [29, Theorem 1]. The converse is not true.
@@
Let $L$ be a small link and $L^{\prime}$ a link such that $L\succeq L^{\prime}$. Suppose, to the contrary, that the link $L^{\prime}$ is not small. By Lemma 5.2, the $3$-orbifold $\mathcal{O}(L^{\prime})$ is not small, that is, it contains an orientable closed incompressible $2$-suborbifold $F$. Therefore, its orbifold fundamental group $G^{\mathrm{orb}}(L^{\prime})$ splits along the orbifold fundamental group $\pi_{1}^{\mathrm{orb}}(F)$ as an amalgamated free product or an HNN extension. In particular, $G^{\mathrm{orb}}(L^{\prime})$ acts non-trivially on the Bass-Serre tree $\mathcal{T}$, without edge inversions, associated to this algebraic decomposition. The epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces a non-trivial action of the group $G^{\mathrm{orb}}(L)$ on the Bass-Serre tree $\mathcal{T}$ without edge inversions. It follows from [65, Corollary 10.2] that the good orbifold $\mathcal{O}(L)$ contains an orientable incompressible $2$-suborbifold, and thus it is not small. By Lemma 5.2, this contradicts the assumption that $L$ is a small link. ∎
@@
(1) A Montesinos knot with $3$ tangles is a small knot by [54, Corollary  4.(a)], hence $K^{\prime}$ must be a small knot by Theorem 1.11(1). Since a small knot is prime, then by Theorem 1.5(3) and Remark 1.6, $K^{\prime}$ is the unknot, a $2$-bridge knot, or a Montesinos knot with $r^{\prime}\leq 3$ tangles, and now $r^{\prime}=3$.
(2) A torus knot is a small knot since the only essential properly embedded surface in its exterior is a fibered annulus. Hence $K^{\prime}$ must be a small knot by Theorem 1.11(1) and then prime. By Theorem 1.5(4), $K^{\prime}$ is the unknot, a $2$-bridge knot, an elliptic Montesinos knot, or a torus knot. Now, $K$ is a $(p,q)$-torus knot with $p$ and $q$ odd, and thus its $2$-fold branched cover $\Sigma_{2}(K)$ is the Brieskorn integral homology sphere $\Sigma(2,p,q)$, see [36, Theorem 7.12]. Hence $K^{\prime}$ cannot be a $2$-bridge knot. If $K^{\prime}$ is an elliptic Montesinos knot, then its $2$-fold branched cover $\Sigma_{2}(K^{\prime})$ is a Seifert fibered integral homology sphere with finite fundamental group. Hence $\Sigma_{2}(K^{\prime})$ must be the Poincaré sphere and $K^{\prime}$ is the $(3,5)$-torus knot (see [8, Affirmation 2.5]). Moreover, if $K^{\prime}$ is a $(p^{\prime},q^{\prime})$-torus knot, then $p^{\prime}$ and $q^{\prime}$ must be odd by [36, Theorem 7.12] since $\Sigma_{2}(K^{\prime})$ is an integral homology sphere. ∎

### v2

# A preorder on the set of links with applications to symmetric unions

@@ Abstract.
For a link $L$ in the $3$-sphere, the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ is defined as a quotient of the link group of $L$. When there exists an epimorphism $G^{\mathrm{orb}}(L)\to G^{\mathrm{orb}}(L^{\prime})$, we denote this by $L\succeq L^{\prime}$ and explore the relationships between the two links. Specifically, we prove that if $L\succeq L^{\prime}$ and $L$ is a Montesinos link with $r$ rational tangles $(r\geq 3)$, then $L^{\prime}$ is either a Montesinos link with at most $r+1$ rational tangles or a certain connected sum. We further show that if $L$ is a small link, then there are only finitely many links $L^{\prime}$ satisfying $L\succeq L^{\prime}$. In contrast, if $L$ has determinant zero, then $L\succeq L^{\prime}$ for every $2$-bridge link $L^{\prime}$. Our main applications concern symmetric unions of knots. In particular, we provide a criterion showing that a given knot does not admit a symmetric union presentation.
@@ 1.1. A preorder defined via orbifolds
Here, the rank $\operatorname{rank}(G)$ of a finitely generated group $G$ is defined to be the minimal number of elements needed to generate $G$.
@@
    If $L$ is a Seifert link whose determinant is non-zero, then $L^{\prime}$ is the unknot, or a $2$-bridge link, or an elliptic Montesinos link or a Seifert link, or a connected sum of $n_{1}$ 22-bridge links and $n_{2}$ elliptic Montesinos links with $n_{1}+2n_{2}\leq\operatorname{rank}\pi_{1}(\Sigma_{2}(L))$.
@@ 1.2. Arborescent links
22-bridge links, Montesinos links, and Seifert links belong to a much wider class of links called arborescent links. A link $L$ is an arborescent link if $\Sigma_{2}(L)$ is a graph manifold, that is to say the geometric decomposition of $\Sigma_{2}(L)$ along essential 2-spheres and tori involves only Seifert fibered pieces. See [14], [37, Chapter 10.7], and [59]. For example, the Kinoshita-Terasaka knot is an arborescent knot. We also get some constraints on a link $\pi$-dominated by an arborescent link.
@@ 1.3. Small links and finiteness
1.4. Symmetric unions
Finally, we use the $\pi$-orbifold group to study a classical construction in knot theory introduced by Kinoshita and Terasaka in [39] and generalized by Lamm [42], which always produces ribbon knots. This construction is a generalization of the connected sum of a knot with its mirror image and called a symmetric union. It gives a $\pi$-domination between the symmetric union and the partial knot (see Section 7). Hence, our results (Theorems 1.5, 1.8 and 1.11) provide strong constraints for a knot to be the partial knot of a symmetric union (see Corollaries 7.3, 7.4, and 7.5). These tools can be applied to the open problem of whether every ribbon knot admits a symmetric union presentation.
The notion of $\pi$-minimal links developed in Section 6 leads to the following criterion for a knot not to be a symmetric union.
Theorem 1.12.
Let $K$ be a $\pi$-minimal knot. If $\operatorname{rank}H_{1}(\Sigma_{2}(K);\mathbb{Z})\geq 3$, then $K$ cannot admit a symmetric union presentation.
We show that the ribbon Montesinos knot $11a_{103}=K(\frac{2}{3},\frac{1}{3},\frac{2}{7})$ is $\pi$-minimal in Proposition 7.13.
@@ Acknowledgments
This study was supported in part by JSPS KAKENHI Grant Numbers JP19K03505, JP23K12974, and JP24H00686. The first and second authors were supported by Soka University International Collaborative Research Grant.
@@ 2. Preliminaries
According to [37, Definition 3.2.2], a link $L\subset S^{3}$ is said to be locally trivial if any $2$-sphere $S$ intersecting $L$ in two points bounds a $3$-ball $B$ such that $B\cap L$ is a trivial arc in $B$. The unknot $U$ is locally trivial. For a split link, it is locally trivial if and only if it is a $2$-component trivial link $U\sqcup U$. A connected sum of two links distinct from the unknot is not locally trivial.
Definition 2.1 ([37, Definition 3.2.4]).
@@
The following factorization theorem will be useful (see [30, Main Theorem], [37, Theorem 3.2.6]).
@@ 2.1. Montesinos links
See [35, Corollary 6.2] for instance. Moreover, $r-2\leq\operatorname{rank}\pi_{1}(\Sigma_{2}(L))\leq r-1$ by [10, Theorem 1.1].
The Seifert fibered manifold $\Sigma_{2}(L)=V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ is determined, up to orientation-preserving homeomorphism, by the rational Euler number $e_{0}\in\mathbb{Q}$ and the set of fractions $\{\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}}\}$ in $\mathbb{Q}/\mathbb{Z}$ up to permutations (see [54], [35, Theorem 1.5]), while the Montesinos link $L(\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ is determined, up to orientation-preserving homeomorphism of $S^{3}$, by the rational Euler number $e_{0}\in\mathbb{Q}$ and the set of fractions $\{\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}}\}$ in $\mathbb{Q}/\mathbb{Z}$ up to dihedral permutations (see [17, Theorem 12.26]). Note that the sign convention for the rational Euler number $e_{0}=\sum_{i}\frac{\beta_{i}}{\alpha_{i}}$ is opposite to the choice made in [52, Section 1].
@@ 2.2. Orbifolds
(1) follows from the proof of the Smith conjecture (see [51]). (2) holds by construction. (3) is proved by the equivariant sphere theorem (see [46, Section 7] or [24]).
@@
Suppose, to the contrary, that $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is centerless. Then $\tilde{\varphi}(Z)=\{1\}$ and $\tilde{\varphi}$ induces an epimorphism from the orbifold group $\Gamma=\pi_{1}(\Sigma_{2}(L))/Z$ onto $\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $\Sigma_{2}(L^{\prime})$ is aspherical, $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is torsion-free. On the other hand, $\Gamma$ is generated by torsion elements. Thus, the image $\tilde{\varphi}(\Gamma)$ must be trivial, and this is impossible. Therefore, $\pi_{1}(\Sigma_{2}(L^{\prime}))$ has a non-trivial center. In particular, it contains an infinite cyclic normal subgroup and it follows from [18, Theorem 1.1] or [27, Corollary 2] that $\Sigma_{2}(L^{\prime})$ admits a Seifert fibration. By the orbifold theorem (see [47]), one can assume that the covering involution $\tau^{\prime}$ on $\Sigma_{2}(L^{\prime})$ is fiber-preserving. If $\tau^{\prime}$ preserves the orientation of the fibers, $L^{\prime}$ is a Seifert link. Otherwise, the underlying space of the base of the Seifert fibration is $S^{2}$ by [48, Section 5.11] and $L^{\prime}$ is a Montesinos link with $r^{\prime}$ rational tangles by [48, Section 2], where
$r^{\prime}\leq\operatorname{rank}\pi_{1}(\Sigma_{2}(L^{\prime}))+2\leq\operatorname{rank}\pi_{1}(\Sigma_{2}(L))+2\leq r+1.$
The center $Z=\langle h\rangle$ of $\pi_{1}(\Sigma_{2}(L))$ is a infinite cyclic normal subgroup in $G^{\mathrm{orb}}(L)$ generated by an element $h$ such that $\bar{\mu}h\bar{\mu}^{-1}=h^{-1}$, where $\bar{\mu}$ is the image of a meridian $\mu$ of $G(L)$ in the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$. This follows from the fact that, for a Montesinos link, the covering involution $\tau$ on $\Sigma_{2}(L)$ is fiber-preserving and reverses the orientation of the fibers (see [48, Section 2] and [12, Theorem 1.3]). The image $\varphi(Z)$ of the center $Z$ of $\pi_{1}(\Sigma_{2}(L))$ is in the center of $\pi_{1}(\Sigma_{2}(L^{\prime}))$.
@@
We now assume that $L^{\prime}$ is not split. Since $L^{\prime}$ is not prime by Theorem 2.2, $L^{\prime}$ decomposes into a connected sum $L^{\prime}=L^{\prime}_{1}\sharp\cdots\sharp L^{\prime}_{n}$ of finitely many prime links which are unique up to permutations ($n\geq 2$). Then, by Lemma 3.1(1)–(3), $\Sigma_{2}(L^{\prime})\cong\Sigma_{2}(L^{\prime}_{1})\sharp\cdots\sharp\Sigma_{2}(L^{\prime}_{n})$ is a non-trivial connected sum of closed, orientable, irreducible $3$-manifolds. In particular, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\cdots\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{n}))$ is a non-trivial free product and hence centerless. The argument given above for the split case shows that each factor $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ must contain some torsion element ($1\leq i\leq n$). Since $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ is the fundamental group of a closed irreducible orientable $3$-manifold for $1\leq i\leq n$, it must be of finite order by [32, Corollary 9.9]. By the orbifold theorem, each $\Sigma_{2}(L^{\prime}_{i})$ is a Seifert $3$-manifold with finite fundamental group, and each link $L^{\prime}_{i}$ is either a $2$-bridge link or an elliptic Montesinos link with three rational tangles. Let $n_{1}$ be the number of 2-bridge factors of $L^{\prime}$ and $n_{2}$ the number of elliptic Montesinos factors of $L^{\prime}$. Then, the number $n=n_{1}+n_{2}$ of prime factors of $L^{\prime}$ verifies $n_{1}+2n_{2}\leq r-1$ since the fundamental group of the $2$-fold branched cover of a $2$-bridge link has rank $1$, that of an elliptic Montesinos link has rank $2$, and that of a Montesinos link with $r\geq 3$ rational tangles has rank at most r1r-1.
@@
By [11], the bridge number $b(K)$ of a Montesinos knot with $r\geq 3$ rational tangles is $r$. Then, the proof follows readily from Theorem 1.5(3) when $K^{\prime}$ is prime. When $K^{\prime}$ is not prime, it is the connected sum of $n_{1}$ 22-bridge knots and $n_{2}$ elliptic Montesinos knots with $n_{1}+2n_{2}\leq r-1$. Schubert’s bridge formula [60] shows that the bridge number minus one is additive by connected sum:
@@
Let $N$ be a closed orientable irreducible $3$-manifold whose JSJ pieces are all hyperbolic. Then $N$ is aspherical and thus $\pi_{1}(N)$ is torsion-free. Moreover, by [34, Theorem VI.1.6(i)], the centralizer in $\pi_{1}(N)$ of any non-trivial element is isomorphic to $\mathbb{Z}$ or $\mathbb{Z}\oplus\mathbb{Z}$ (see also [26]), hence abelian. Then, Theorem 1.8 follows from Theorem 4.7 and Corollary 4.10, where the proof of Theorem 4.7 is based on Lemmas 4.1 and 4.3 below.
@@
Since $M_{0}$ is a rational homology solid torus, one can choose a simple closed curve $\alpha\subset\partial M_{0}$ such that the Dehn filling $M_{0}(\alpha)$ is a rational homology sphere (see [62, Lemma 3.2] for example). The natural epimorphism
@@
Combining Theorem 4.7 with [34, Theorem VI.1.6(i)], one has the next corollary.
@@ 5. Small links and finiteness
If $L$ is not a split link, $E(L)$ is irreducible and a properly embedded essential surface $F$ in $E(L)$ is boundary-incompressible, see [53, Lemma 2.1]. If $\Sigma_{2}(L)$ does not contain any incompressible surface, then the link $L$ is small by [28, Theorem 1]. The converse is not true.
@@
Let $L$ be a small link and $L^{\prime}$ a link such that $L\succeq L^{\prime}$. Suppose, to the contrary, that the link $L^{\prime}$ is not small. By Lemma 5.2, the $3$-orbifold $\mathcal{O}(L^{\prime})$ is not small, that is, it contains an orientable closed incompressible $2$-suborbifold $F$. Therefore, its orbifold fundamental group $G^{\mathrm{orb}}(L^{\prime})$ splits along the orbifold fundamental group $\pi_{1}^{\mathrm{orb}}(F)$ as an amalgamated free product or an HNN extension. In particular, $G^{\mathrm{orb}}(L^{\prime})$ acts non-trivially on the Bass-Serre tree $\mathcal{T}$, without edge inversions, associated to this algebraic decomposition. The epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces a non-trivial action of the group $G^{\mathrm{orb}}(L)$ on the Bass-Serre tree $\mathcal{T}$ without edge inversions. It follows from [64, Corollary 10.2] that the good orbifold $\mathcal{O}(L)$ contains an orientable incompressible $2$-suborbifold, and thus it is not small. By Lemma 5.2, this contradicts the assumption that $L$ is a small link. ∎
@@
(1) A Montesinos knot with three rational tangles is a small knot by [53, Corollary  4.(a)], hence $K^{\prime}$ must be a small knot by Theorem 1.11(1). Since a small knot is prime, then by Theorem 1.5(3) and Remark 1.6, $K^{\prime}$ is the unknot, a $2$-bridge knot, or a Montesinos knot with $r^{\prime}\leq 3$ rational tangles, and now $r^{\prime}=3$.
(2) A torus knot is a small knot since the only essential properly embedded surface in its exterior is a fibered annulus. Hence $K^{\prime}$ must be a small knot by Theorem 1.11(1) and then prime. By Theorem 1.5(4), $K^{\prime}$ is the unknot, a $2$-bridge knot, an elliptic Montesinos knot, or a torus knot. Now, $K$ is a $(p,q)$-torus knot with $p$ and $q$ odd, and thus its $2$-fold branched cover $\Sigma_{2}(K)$ is the Brieskorn integral homology sphere $\Sigma(2,p,q)$, see [35, Theorem 7.12]. Hence $K^{\prime}$ cannot be a $2$-bridge knot. If $K^{\prime}$ is an elliptic Montesinos knot, then its $2$-fold branched cover $\Sigma_{2}(K^{\prime})$ is a Seifert fibered integral homology sphere with finite fundamental group. Hence $\Sigma_{2}(K^{\prime})$ must be the Poincaré sphere and $K^{\prime}$ is the $(3,5)$-torus knot (see [8, Affirmation 2.5]). Moreover, if $K^{\prime}$ is a $(p^{\prime},q^{\prime})$-torus knot, then $p^{\prime}$ and $q^{\prime}$ must be odd by [35, Theorem 7.12] since $\Sigma_{2}(K^{\prime})$ is an integral homology sphere. ∎